# PromptRPA: Generating Robotic Process Automation on Smartphones from Textual Prompts

### v1

@@ 8. Conclusion
This paper presents PromptRPA as an effective solution in Robotic Process Automation, particularly addressing the barriers of technical complexity that deter its mass adoption. Utilizing the multi-agent framework and supplementing online tutorials, PromptRPA is capable of interpreting unrestricted textual prompts to perform a wide range of RPA tasks. Performance evaluations showcase a marked rise in success rates, ascending from an initial 22.28% to a remarkable 95.21%, with an average of only 1.66 user interventions per new task. Even without user interventions, a success rate of 83.68% is achieved. User studies further validate the effectiveness of PromptRPA in delivering a user-friendly and intelligent experience. Beyond demonstrating the viability of text-driven control in RPA, our research paves the way for continued exploration, emphasizing enhanced functionality and versatility. Upcoming efforts aim to minimize the burden of user interaction, engage a wider user group, and boost the system’s proficiency in managing complex tasks.
@@ Appendix A Mult-Agent Design
A.1. Grounding Agent
In this section, we provide detailed descriptions of the strategies employed by the Grounding Agent. Initially, based on the current page’s layout hierarchy, the agent identifies and selects all widgets that either possess text or have interactive attributes, thereby constructing a set of potential operation candidates. Subsequently, the agent employs the following strategies to predict the operational widget within these candidates:
@@ A.1. Grounding Agent
    Historical RPA invocation: If the instruction has been previously mapped to an operation, the agent has the capability to refer to stored widget data in the historical RPA repository. It enables the agent to direct the operation to a corresponding widget on the current interface. To evaluate the similarity between the candidate widget and the stored widget, we employ a heuristic rule-based similarity function, as described in the work of Huang et al. (Huang et al., 2023). This function considers 26 features of a widget, including its style and position in images, its style and position in layouts, and its role in page segmentation. The similarity thresholds have been established using Receiver Operating Characteristic (ROC) curve analysis.
@@
      Strict text matching: It involves searching for a widget on the page that closely matches the object description ($Instruction.Object$). Similarity is defined as Equation 1, where $Widget.text$ denotes the text and content description contained in the widget’s layout information, and $LCS$ signifies the length of the longest common sub-string between two strings. A match is confirmed only if the similarity exceeds an empirically preset threshold.
      (1) $Sim(Widget,Instruction.Object)$ $=\frac{2\times LCS(Widget.text,Instruction.Object.text)}{length(Widget.text)+length(Instruction.Object.text)}$
@@
      Textual semantic matching: This approach leverages the operation and object description from the instruction. The agent enumerates potential widgets with relevant information, including their text, content description, position on the current page, and interactive attributes such as clickability and editability. Subsequently, the LLM considers each widget’s semantic relevance to the instruction and then predicts the most suitable widget for executing the given instruction.
    3. (c)
      Exploratory mechanism: In instances where the operation widget proves difficult to identify, this mechanism becomes particularly instrumental. The agent presents the LLM with a wealth of contextual information to refer to:
      1. (i)
        The function description in the model.
      2. (ii)
        The system’s current reference instructions, upcoming potential instructions, and completed instructions.
      3. (iii)
        Detailed attributes of all interactive widgets, such as their text, position on the current page, and interactive attributes.
      4. (iv)
        Descriptions of future pages each interactive widget may lead to. If these pages’ layout and screenshot data have been stored in the mobile interaction graph, the Mobile Semantics Agent can generate textual descriptions for them.
      5. (v)
        The reason for the previous prediction that the agent responded.
      6. (vi)
        The evaluation made by the Assessment Agent on the previous prediction.
A.2. Assessment Agent
The Assessment Agent utilizes the LLM with a rich set of contextual information, encompassing:
  • Function Description: It outlines the context or purpose of the current task.
  • Current Page Description: Sourced from the Mobile Semantics Agent, it provides context about the mobile GUI that the agent is interacting with.
  • Operation History: A log of operations that the agent has previously performed, useful for understanding the context and making future predictions.
  • Current and Future Instructions: The immediate instruction the agent is focusing on, as well as potential future instructions that might influence the current operation. Although these instructions may not be strictly adhered to, they serve as valuable reference points.
  • Details about the Predicted Operation: In addition to information about the widgets predicted by the Grounding Agent, it also incorporates reasoning information from the Grounding Agent:
    • The rationale behind the Grounding Agent’s prediction, explaining why this operation was chosen.
    • Potential effects of the predicted operation, such as future pages the operation may lead to, which might be found in the mobile interaction graph, to facilitate reasoned decisions.
  • History of Attempts: The agent keeps a record of previously attempted and failed interactions with pages and widgets. This is designed to prevent dead loops during the exploration process.
The LLM is then tasked with choosing between the current predicted operation and a possible rollback, also furnishing a confidence score for its decision. The confidence score is used to determine whether human intervention is necessary. An empirical threshold is set, below which the system automatically prompts the user for confirmation.
The agent has several built-in guidelines for pathway exploration, to effectively oversee the decision-making of the Grounding Agent:
  • Limit Scrolling: The agent minimizes excessive scrolling on a page to avoid worthless exploration.
  • Rationality Check: When considering the addition of new steps, such as clicking a widget to explore tasks, the agent evaluates their logical solidity.
  • Outcome Validation: Should the outcome of a previous step fail to meet expectations or prove irrelevant to the RPA task, the agent opts for the ‘RETRACT’ option.
  • Task Completion: If the agent determines that the task is complete, it selects the ‘FINISH’ option, discarding unnecessary steps.
@@ Appendix B Accumulated Knowledge
B.1. Construction of the Historical RPA Repository
The Historical RPA Repository serves as an archive for successfully executed RPA tasks. Since the Grounding Agent doesn’t strictly adhere to pre-generated instructions, the repository focuses on storing the actual sequences of operations executed. This subsection details how these operation sequences are reverse-engineered into corresponding instructions and step descriptions.
@@ B.1. Construction of the Historical RPA Repository
    Operation Sequence Storage: The first step is to archive the actual sequence of operations executed by the Grounding Agent, which is the most reliable record of the RPA task. The system archives all layout hierarchical information and screenshots of the pages visited. It also logs the operation object’s index in the page’s layout tree to retrieve the object’s layout information. Additionally, since the object’s position is recorded, it allows for capturing its corresponding image on the full-page screenshot.
@@
    Reverse Engineering to Instructions: Using the stored operation sequences, the system reverse-engineers to produce the corresponding instructions. Each operation is traced back to the potential instruction that could have prompted it, adhering to the format specified in Table 4.1.3. In this context, the operation types and parameters have already been logged by the system during the triggering of the operation. The descriptions of operation objects are derived from their texts, content descriptions, positions, and the average RGB values of their respective images.
@@
    Generation of Step Descriptions: Alongside instructions, the system also formulates step descriptions. They are rephrased versions of the instructions, articulated in natural language. While these historical step descriptions may not directly contribute to the system’s operational efficiency, they enhance the user experience by making the automated processes more transparent and easier to understand.
@@ C.1. Objective Results
Manual First Time Using PromptRPA Second Time Using PromptRPA

### v2

@@ 8. Conclusion
This paper presents Prompt2Task as an effective solution in UI task automation, particularly addressing the barriers of technical complexity, thereby facilitating its broader application. Utilizing the multi-agent framework and supplementing online tutorials, Prompt2Task is capable of interpreting unrestricted textual prompts to perform a wide range of automation tasks. As knowledge is accumulated from user usage and feedback, Prompt2Task improves task success rates and efficiency. Performance evaluations showcase a marked rise in success rates, ascending from an initial 22.28% to a remarkable 95.24%, with an average of only 0.69 user interventions per new task. Even without user interventions, a success rate of 78.16% is achieved. User studies further validate the effectiveness of Prompt2Task in delivering a user-friendly and intelligent experience. Upcoming efforts aim to minimize the burden of user interaction, engage a wider user group, and boost the system’s proficiency in managing complex tasks.
Acknowledgements.
This work is supported by the Natural Science Foundation of China under Grant No. 62132010, Beijing Key Lab of Networked Multimedia, Institute for Artificial Intelligence, Tsinghua University (THUAI), Beijing National Research Center for Information Science and Technology (BNRist), 2025 Key Technological Innovation Program of Ningbo City under Grant No. 2022Z080, Beijing Municipal Science and Technology Commission, Administrative Commission of Zhongguancun Science Park No.Z221100006722018, and Science and Technology Innovation Key R&D Program of Chongqing.
@@ Appendix A Mult-Agent Design
The following provides the technical and prompt details for each agent.
A.1. Analysis Agent
The agent processes the input through GPT-4 to produce a function description and step description as reflected in the input, with prompt details shown in Figure 17.
Figure 17. Prompts used by the Analysis Agent.
A.2. Retrieval Agent
The agent uses GPT-4 to extract the step description from the raw text of search results, with prompt details shown in Figure 18.
Figure 18. Prompts used by the Retrieval Agent.
A.3. Parsing Agent
The agent uses GPT-4 to accomplish the transformation from step descriptions in natural language to formalized instructions, with prompt details shown in Figure 19.
Figure 19. Prompts used by the Parsing Agent.
A.4. Grounding Agent
In this section, we provide detailed descriptions of the strategies employed by the Grounding Agent. Initially, based on the current page’s layout hierarchy, the agent identifies and selects all widgets that either possess text descriptions or have interactive attributes, thereby constructing a set of potential operation candidates. Subsequently, the agent employs the following strategies to predict the operational widget within these candidates:
@@ A.4. Grounding Agent
    Historical task invocation: If the instruction has been previously mapped to an operation, the agent can refer to stored widget data in the historical task repository. It enables the agent to direct the operation to a corresponding widget on the current interface. To evaluate the similarity between the candidate widget and the stored widget, we employ a heuristic rule-based similarity function, as described in the work of Huang et al. (Huang et al., 2023). This function considers 26 features of a widget, including its style and position in images, its style and position in layouts, and its role in page segmentation. The similarity thresholds have been established using Receiver Operating Characteristic (ROC) curve analysis.
@@
[…]
      (2) $Sim(\text{Widget},\text{Instruction}.\text{Object})$ $=\frac{2\times LCS(\text{Widget}.\text{text},\text{Instruction}.\text{Object}.\text{text})}{length(\text{Widget}.\text{text})+length(\text{Instruction}.\text{Object}.\text{text})}$
@@
      Redirect, Add, Skip, Expand, Block: The agent enumerates potential widgets with relevant information, including their text, content description, position on the current page, interactive attributes and so on. Subsequently, the LLM considers the current context to choose a strategy and then predicts the most suitable widget for executing the operation. Details of the prompt are illustrated in Figure 20.
Figure 20. Prompts used by the Grounding Agent.
A.5. Assessment Agent
The Assessment Agent utilizes the LLM with a rich set of contextual information to make decisions: follow, change, retract, or finish, while also describing if it is confident for its decision, as shown in Figure 21. If the agent is unconfident in its decision, the system automatically prompts the user for assistance.
Figure 21. Prompts used by the Assessment Agent.
@@ Appendix B Accumulated Knowledge
B.1. Construction of the Historical Task Repository
The Historical Task Repository serves as an archive for successfully executed automation tasks. Since the Grounding Agent doesn’t strictly adhere to pre-generated instructions, the repository focuses on storing the actual sequences of operations executed. This subsection details how these operation are reverse-engineered into corresponding instructions and step descriptions.
@@ B.1. Construction of the Historical Task Repository
    Operation Sequence Storage: The first step is to archive the actual sequence of operations executed by the Grounding Agent, which is the most reliable record of the automation task. The system archives all layout hierarchical information and screenshots of the pages visited. It also logs the operation object’s index in the page’s layout tree, which can be used to retrieve the object’s layout information. Additionally, since the object’s position is recorded, it allows for capturing its corresponding image on the full-page screenshot.
@@
    Reverse Engineering to Instructions: Using the stored operation sequence, the system reverse-engineers to produce the corresponding instructions. Each operation is mapped to the potential instruction that could have prompted it, adhering to the format specified in Table 4.1.3. In this context, the operation types and parameters have already been logged by the system during the triggering of the operation. The descriptions of operation objects are derived from their texts, content descriptions, positions, and the average RGB values of their corresponding images.
@@
    Generation of Step Descriptions: Alongside instructions, the system also formulates step descriptions. They are rephrased versions of the instructions, articulated in natural language, generated by GPT-4.
@@ B.2. Construction of the Context Library
Here is an example of records in the context library. Each time a task is successfully completed, the context library integrates new information using GPT-4, with prompt detailed in Figure 22.
    "hardwareInfo": {
        "deviceName": "mySmartPhone",
        "phoneModel": "HUAWEI Mate30",
        "battery": "0.23",
        "availableWifi": ["Network1", "Network2"]
    "systemSetting": {
        "display": {
            "darkmode": true,
            "autoBrightness": true
    "privacyAndSecurity": {
        "locationAllowedServices": {
            "allowed": ["Maps", "WeChat", "Lark"]
        "permissions": {
            "camera": true,
            "location": true,
            "contacts": false
    "usingApp": ["App1", "App2"],
    "operatingSystem": "HarmonyOS 2.0",
    "location": "Beijing, China",
    "timeOfDay": "2023-04-12T14:30:00Z",
    "networkConnectivity": "Wi-Fi",
    "userPreferences": {
        "preferredLanguage": "en",
        "themeSettings": "dark"
    "connectedDevices": {
        "bluetoothHeadphones": "HUAWEI Freebuds",
        "smartwatch": "HUAWEI Watch"
Figure 22. Prompts used to update the context library based on the currently completed automation task.
B.3. Typical Cases
This section illustrates typical cases where accumulated knowledge optimizes the system in various ways, while also explaining how the system handles outdated knowledge.
B.3.1. Case 1
Case 1 demonstrates the optimization of the Grounding Agent by the historical task repository.
Suppose the current instruction is ¡type: click, object: “Account”¿. As shown in Figure 23, the current UI suggests that “Me” should be clicked based on semantic similarity. During the initial execution, the Grounding Agent uses the LLM and chooses the redirect strategy to locate the target widget. This process is relatively slow and may result in inconsistent outcomes. If the target widget cannot be found, the Grounding Agent will choose to block, prompting the user to demonstrate the correct operation.
After the first execution, in subsequent executions, the system leverages the Prompt2Task model stored in the historical task repository. Knowing that this step is clicking “Me”, the Grounding Agent can directly locate the target widget using heuristic rules, resulting in faster and more reliable outcomes.
If the app undergoes significant updates such that “Me” no longer exists or the widget style changes drastically, making it difficult to locate the widget using heuristic rules, the Grounding Agent will refer to the instruction ¡type: click, object: “Me”¿ and use the LLM to infer the operation again (using strategies such as redirect, add, skip, expand, or block). Even if the knowledge in the historical task repository is outdated, it still provides guidance by including more target details in the prompt, as shown in the prompt engineering of the Grounding Agent in Appendix A.4. After completing the new task, the historical task repository will store the updated target details.
Figure 23. Case 1: The instruction indicates clicking “Account”, but the current UI suggests that “Me” should be clicked instead.
B.3.2. Case 2
Case 2 demonstrates the optimization of the mobile semantics module in the Grounding Agent by the historical task repository, which then enhances the agent’s prediction.
Suppose the current instruction is ¡type: click, object: “more”¿, and the current UI is as shown in Figure 24. During the initial execution, the agent is unable to identify “more” due to insufficient UI information (no text labels and pixel-based recognition failed). Therefore, the Grounding Agent enters a block state and requests the user to demonstrate the correct operation.
After the first execution, in subsequent executions, the mobile semantics module learns from the Prompt2Task model stored in the historical task repository and knows that the icon image corresponds to the “more” label. Therefore, it can provide this information to the Grounding Agent, which can then use strict text matching to directly locate and click the widget.
As long as the icon image does not change significantly, the mobile semantics module can label the icon as “more”. Otherwise, the process of requiring user intervention to identify the icon will be repeated, and after task completion, the historical task repository will be updated accordingly.
Figure 24. Case 2: The instruction indicates clicking “more”, but the icon’s information is insufficient for the agent to locate it, likely requiring user demonstration upon first encounter.
B.3.3. Case 3
Case 3 demonstrates the optimization of the Analysis Agent by the context library.
Figure 25. Case 3: An example of auto-completing the textual prompt.
As shown in Figure 25, the user input is “How to adjust the sound quality of earphones”. During the initial encounter, the Analysis Agent needs to query the user for the earphones’ model. Once the user provides the information, the context library integrates it, as described in Appendix B.2.
In subsequent instances of similar tasks, the agent will proactively ask the user if he/she needs to “adjust the sound quality of HUAWEI Freebuds Pro”, requiring only a simple confirmation from the user.
If the user’s context changes, such as a different earphone model, the context library may provide incorrect information. In such cases, the user can reject the agent’s recommendation, and the agent will engage in a dialogue with the user to correct and then update the context library.
B.3.4. Case 4
Case 4 demonstrates the optimization of the Parsing Agent by the instruction set.
Suppose the current step description is “copy the title”, which is a complex instruction that should be decomposed into two instructions:
  • ¡type: longclick, object: title widget¿ (Figure 26)
  • ¡type: click, object: “copy”¿ (Figure 26)
Figure 26. Case 4: The instruction is a complex copy command that requires first long-pressing the title and then selecting “copy” from the options that pop up.
During the first execution, since it is not included in the basic instruction definition, the Parsing Agent can only generate ¡type: copy, object: “title”¿. Consequently, the Grounding Agent cannot match the corresponding operation and must rely on the “add” strategy or user demonstration to successfully execute the task. After task completion, the user maps the original compound instruction to the generated decomposed instructions, and the instruction set stores this correspondence for future use.
In subsequent executions, the Parsing Agent will automatically decompose the compound instruction by finding similar examples in the instruction set. Then it will allow the Grounding Agent to potentially find the target widget directly through heuristic rules. Even if it cannot, it will reduce the difficulty of LLM reasoning in the Grounding Agent.
If the app undergoes significant updates, changing the execution steps of this compound instruction, the instructions generated by the Parsing Agent may be incorrect. Consequently, the Grounding Agent will rely on the LLM to redeploy strategies (redirect, add, skip, expand). If it encounters a block, it will require user demonstration.
B.3.5. Case 5
Case 5 demonstrates the optimization of the Grounding Agent and the Assessment Agent by the mobile interaction graph.
Suppose the current step description is “share the link with a friend”, and the current UI is shown in Figure 27. There are four interactive widgets available for operation from left to right. Suppose the mobile interaction graph already contains the resulting pages after triggering these widgets (with the mobile semantics module in the Grounding Agent generating textual descriptions for each page):
  • 0 $\rightarrow$ search result page
  • 1 $\rightarrow$ forwarding list
  • 2 $\rightarrow$ cart popup
  • 3 $\rightarrow$ tools page
As shown in Figure 20 and Figure 21, this information is incorporated into the prompts for the Grounding Agent and the Assessment Agent, aiding them in making accurate predictions and judgments.
If a significant app update alters the page transition relationships, the mobile interaction graph might provide incorrect information, potentially leading both agents to erroneous decisions. If multiple erroneous attempts occur, the Assessment Agent will request user intervention.
Figure 27. Case 5: The current UI contains 4 candidate widgets. The agents will consider various information for each widget, including the triggered future pages, to make the final decision.
@@ C.1. Objective Results
Manual Prompt2Task (First Time) Prompt2Task (Second Time)